The number of pounds of tomatoes y consumed per person in the US in year x is given by 5x-22y+286=0,
where x=0 corrsponds to 1980. Find the year in which tomato consumption per person is 18lb.

Answer by jorel1380(3719)

5x-22y+286=0
5x=22y-286
22(18)-286/5=x
x=22
1980+22= The year 2002..